OpenStudy (anonymous):

5y + -7 + y = -1(y + 3y

OpenStudy (anonymous):

Get y by itself. Combine like terms. Start off with adding -7 to both sides.

OpenStudy (anonymous):

So you would have 5y+y=-8(y+3y) Like I said, combine like terms. (y) 6y = -8 (4y) Can you do it from there, or do you want me to walk you through it?

OpenStudy (anonymous):

walk me through , i'm very confused. I have a very hard time with solving equation

OpenStudy (anonymous):

6y=-8(4y) Multiply -8 by 4y. 6y = -32y Divide both sides by 6. y = 7/10